4.00x + 6.00y = 2340
x + y = 480

Answer by stanbon(75887)

4.00x + 6.00y = 2340
x + y = 480
----------------------
4x + 6y = 2340
x + y = 480
--------------------
2x + 3y = 1170
x + y = 480
------------------------
Multiply thru the 2nd equation by 2:
2x + 3y = 1170
2x + 2y = 960
-------------------
Subtract 2nd from 1st:
y = 210
-----
substitute into x + y = 480 to solve for "x":
x + 210 = 480
x = 270
==================
Cheers,
Stan H.